OpenStudy (anonymous):

Describe the graph of the following inequality: y<-3/2+4 What are the x and y intercepts of the line that separates the domains? Is the line a solid line or a dashed line? Are the points that satisfy this inequality above or below the line?

OpenStudy (anonymous):

it is dashed line because it's not equal to. Well -3/2 is your slope and you start at your 4 and it is shaded below that because it is a less than sign.

OpenStudy (anonymous):

Any idea what the x and y intercepts would be?

OpenStudy (anonymous):

is your inequality y < -3/2x + 4 ?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

if so, to find your y intercept, in y = mx + b form, the b is your y intercept. Another way to find the y intercept is to sub in 0 for x. To find your x intercept, sub in 0 for y. y intercept is : y < -3/2x + 4 y < -3/2(0) + 4 y < 4 -- this is your y intercept. x intercept is : y = -3/2x + 4 0 = -3/2x + 4 3/2x = 4 x = 4/(3/2) x = 4 * 2/3 x = 8/3 or if it makes it easier for you to graph, it is 2 2/3 It will be a dashed line and the points that satisfy this line will be below the line

OpenStudy (anonymous):

So if I had 4y >(or less than) 12x-20 would it be the same thing?

OpenStudy (anonymous):

what do you mean, " would be the same thing ?"

OpenStudy (anonymous):

Would you use the same type formula to solve this?

OpenStudy (anonymous):

you want to put your inequality in y = mx + b form yo find the slope. But as far as finding the x and y intercept, just sub in 0 for x or y, depending, of course, which one you are looking for.

OpenStudy (anonymous):

4y > 12x - 20 -- divide both sides by 4 y > 3x - 5---slope is 3, y intercept is -5 x intercept is : 4(0) > 12x - 20 -12x > -20 x < 20/12 = 1 2/3

OpenStudy (anonymous):

Any idea if this would be a solid or dashed line?

OpenStudy (anonymous):

if there is no equal sign involved, then it is a dashed line. When an equal sign is involved, then it is a solid line.